a rectangular garden has the width of 5 feet and a length of 10 feet. if an equal amount is added to both the width and length, the area is inceased to 104 square feet. what is this amount?

Answer by jorel1380(2518)

(5+x)(10+x)=104
(x+5)(x+10)=104
x2+15x+50-104=0
x2+15x-54=0
(x+18)(x-3)=0
x=-18 or 3
Throwing out the negative answer, we get the amount added to be 3ft..